# Zewei Zhou Tianhui Cai11footnotemark: 1 Seth Z. Zhao  Yun Zhang  Zhiyu Huang Bolei Zhou  Jiaqi Ma

### v1

@@ 1 Introduction
Autonomous driving systems typically adopt a modular paradigm, decomposing the driving task into different sub-modules, such as perception [1, 2, 3], prediction [4, 5, 6], and planning [7, 8, 9]. While this design enables structured development, it may cause error accumulation and a lack of joint optimization across modules, leading to suboptimal performance [10, 11]. End-to-end autonomous driving has gained prominence with a unified model architecture that maps raw sensor inputs directly to final driving actions. These models are trained on human driving data, enhancing scalability and human-like behavior. Vision-based approaches have garnered significant interest due to their affordability and ease of deployment [12, 13, 14, 15].
@@
However, conventional end-to-end methods [16, 17, 18, 19] primarily focus on imitating expert trajectories, lacking essential world knowledge for understanding and reasoning about surrounding environments, particularly in long-tail or challenging scenarios. Recent advances in Vision-Language Models (VLMs) [20, 21, 22] have gained significant interest by introducing models capable of leveraging extensive world knowledge and powerful reasoning. These models have shown strong potential in improving adaptability and scalability across diverse driving scenarios [23, 24, 25, 26, 27, 28, 29]. Building upon VLMs, Vision-Language-Action (VLA) models extend this capability to action generation, enabling embodied agents, such as robots [30, 31, 32] and autonomous vehicles [33, 34], to produce feasible physical actions based on visual observations and language instructions.
@@
Despite recent progress, existing VLA models face two critical limitations in autonomous driving, as illustrated in Fig. 2. 1) Physically-infeasible or complex structure for action generation. Some models generate textual actions or waypoints directly using VLMs [35, 36, 37], but these outputs can be physically infeasible and suffer from mode collapse. To address this, recent approaches introduce intermediate meta-actions [38, 39, 40] or latent action tokens [41, 42, 43], which are then processed by downstream planners or decoders to produce physically feasible trajectories. However, the intermediate representations either break the end-to-end optimization paradigm or increase model complexity and training overhead. 2) Inflexible and inefficient reasoning across diverse scenarios. Most existing models [44, 45] employ a fixed reasoning strategy, lacking the ability to adaptively switch between direct action outputs for straightforward scenarios and chain-of-thought (CoT) reasoning for complex ones. Although DriveVLM [46] introduces a dual-process paradigm, it relies on separate modules (i.e., a VLM for slow reasoning and a conventional end-to-end model for fast responses), which results in a complicated architecture, increased training overhead, and limited scalability [47].
@@
To overcome these limitations, we propose AutoVLA, an end-to-end autonomous driving framework that directly integrates physical action tokens into a pretrained VLM backbone, enabling direct learning of an autoregressive planning policy, as illustrated in Fig. 1. Our unified architecture seamlessly integrates reasoning and action generation, allowing adaptive switching between direct trajectory generation and CoT reasoning. In supervised fine-tuning (SFT), we leverage both trajectory-only data and CoT reasoning data to equip the model with dual-process capabilities (fast and slow thinking). Furthermore, we propose reinforcement fine-tuning (RFT) [48], utilizing Group Relative Policy Optimization (GRPO) [49] with verifiable planning reward functions. This enables adaptive reasoning that balances planning accuracy and efficiency. The RFT method not only improves planning performance but also runtime efficiency by minimizing unnecessary reasoning.
We extensively evaluate AutoVLA using real-world datasets, including nuPlan [50, 51], Waymo [52], nuScenes [53], and simulation datasets such as CARLA [54, 55]. Experimental results demonstrate that AutoVLA achieves superior performance across various end-to-end autonomous driving benchmarks under both open-loop and closed-loop tests. Empirical results validate that our RFT approach markedly improves planning performance, enables adaptive fast and slow thinking capabilities, and reduces runtime by minimizing redundant reasoning. The main contributions of this paper are summarized as follows:
@@ 2 Related Work
End-to-end Autonomous Driving. End-to-end autonomous driving approaches have made significant advances in recent years [10, 11, 56, 57, 58, 59, 60, 61, 62, 63, 64]. Methods such as UniAD [65] and VAD [66] explicitly integrate multiple driving tasks from perception to planning in a unified Transformer architecture, thereby enhancing planning performance. ParaDrive [67] discusses the necessary components within end-to-end driving architectures. Additionally, GenAD [68] and DiffusionDrive [69] adopt generative models to maintain trajectory continuity and produce multi-modal driving trajectories. However, integrating world knowledge into end-to-end driving systems remains challenging due to bottlenecks in semantic reasoning [34] and limited adaptability in complex environments [70].
@@
VLA and VLM for Autonomous Driving. The gap between semantic reasoning and physical actions remains a critical challenge for VLA and VLM in end-to-end autonomous driving. Current research broadly follows three directions. The first directly formulates driving as a language-centric problem, utilizing VLMs for scenario understanding through caption generation [71, 72, 73] or question answering [74, 75]. The second direction leverages VLA or VLM to produce high-level meta-actions or driving decisions [17, 38, 39, 40], which are used to either supervise [12, 76, 77, 78] or guide [39, 79] traditional planners or end-to-end models. Although these approaches facilitate integration, they prevent full end-to-end optimization. Thus, a third direction directly integrates VLMs with action generation into VLA models, enabling the direct prediction of latent action tokens [34, 35, 36, 43] or final driving trajectories [37, 44, 80, 81, 82, 83]. However, simple trajectory decoders employed in these methods (e.g., MLP [41, 84] or GRU [45]) may produce impractical trajectories and suffer from modal collapse. To address this issue, ORION [42] incorporates generative planners into VLM architectures, enhancing trajectory feasibility but increasing model complexity and computational demands. In our work, we integrate a physical action codebook for vehicle motion into a pretrained VLM to effectively bridge the semantic reasoning and physical action space.
@@
Reinforcement Fine-tuning. RFT [48] has shown considerable promise in enhancing the performance and adaptability of LLMs, as demonstrated in DeepSeek-R1 [22]. In autonomous driving, Gen-Drive [7] and TrajHF [85] employed the RFT to align the trajectory generation model with safety constraints and human driving preferences. RAD [86] combined 3D Gaussian splatting to generate scenarios and conduct closed-loop RL training. However, the application of RFT in end-to-end VLM/VLA-based autonomous driving remains nascent. While previous methods, such as AlphaDrive [38], utilize GRPO instead of direct preference optimization (DPO) [87] to enhance planning performance and ensure training efficiency and stability, they are still limited to simplified settings involving only high-level meta-actions. In this work, we advance this direction by applying RFT to optimize the end-to-end VLA framework in both scene reasoning and low-level planning, and we adopt GRPO to accelerate convergence and enhance training stability.
@@ 3 AutoVLA
The proposed AutoVLA framework consists of two main components, as shown in Fig. 1. 1) VLM Backbone: It is capable of processing visual and textual input and generating corresponding tokens (reasoning and action), employing a unified autoregressive Transformer decoder. 2) Physical Action Token Generation: We extend the language model decoder to output physical action tokens that directly correspond to vehicle movements. These tokens are designed to comply with physical constraints and can be reliably translated into physically feasible planning trajectories.
Training of AutoVLA is conducted in two stages, as illustrated in Fig. 3. 1) Supervised Fine-Tuning uses ground-truth trajectory data and distills high-quality reasoning data from a large-scale VLM. 2) Reinforcement Fine-Tuning uses task-specific reward functions to optimize planning performance while improving the running efficiency by minimizing unnecessary reasoning. The details of our model and training process are illustrated below.
@@ 3.1 Framework
Base VLM Model. We adopt Qwen2.5-VL-3B [21] as the vision-language backbone of AutoVLA. Qwen2.5-VL is a series of powerful multimodal large language models that possess strong visual understanding capabilities, and the open-source nature of the Qwen2.5-VL model facilitates task-specific fine-tuning. The 3B variant offers a good trade-off between efficiency and performance, making it suitable for deployment in onboard devices.
Action Tokenization. To enable trajectory planning within the language model, we discretize continuous vehicle trajectories $\mathbf{P}\in\mathbb{R}^{\tau\times d}$ into a sequence of physical action tokens $\mathbf{a}=[a_{1},\dots,a_{T}]$, where $a_{t}\in\mathcal{A}$, $T$ is the length of the tokenized predicted trajectory and each token is represented by short-term spatial position and heading movement $(\Delta x,\Delta y,\Delta\theta)$. This transforms the planning task into a next-token prediction problem, which can be conducted within the language model. We build our action codebook $\mathcal{A}=\{a_{1},a_{2},\dots,a_{K}\}$ using a K-disk clustering method [88, 89, 90], which covers the majority of vehicle movement patterns. Finally, we obtain a vehicle motion codebook that consists of $K=2048$ discrete action tokens. Following [30, 91], these action tokens are incorporated into the VLM as additional tokens (i.e., <action_0>, <action_1>, …). During inference, the model outputs a sequence of these action tokens, which are subsequently decoded into a planning trajectory using the action codebook. More details about action tokenization and trajectory decoding are provided in the Supplementary Material.
Unified Reasoning and Action. AutoVLA unifies reasoning and action generation within a single autoregressive Transformer framework, enabling adaptive switching between fast and slow thinking depending on the driving scenario. In fast thinking mode, AutoVLA directly predicts physical action tokens without generating long reasoning steps, enabling rapid responses in straightforward scenarios. In contrast, slow thinking mode involves structured CoT reasoning, where the model first analyzes the environment, identifies critical elements, and reasons through potential outcomes before deciding on the final driving action. To enable this dual thinking capability, AutoVLA is trained with a mixture of direct action supervision and reasoning-augmented data. We design system prompts and response formats to support both modes consistently.
@@ 3.2 Reasoning Data
Reasoning data provides high-quality CoT annotations that are essential for training VLMs with reasoning capabilities [42]. In driving tasks, reasoning involves understanding complex semantics and interactions in dynamic environments [92, 93, 94, 95]. Despite its importance, the development of a high-quality, large-scale driving reasoning dataset remains a key challenge due to three major limitations: 1) limited scenario diversity and repetitive examples, 2) inadequate representation of critical perceptual cues, such as traffic signs and vehicle indicator signals, 3) low-quality reasoning process, such as repeatedly stopping at a stop sign without justification.
To address these issues, we propose an automated reasoning annotation pipeline using the advanced Qwen2.5-VL-72B model [21]. This pipeline enables automatic generation of high-accuracy reasoning annotations and supports knowledge distillation from a large capable model to a more compact target model. The pipeline generates structured reasoning annotations across four key components: detailed scene descriptions, identification of crucial objects, prediction of surrounding agents’ intentions, and determination of appropriate driving actions. To regulate the reasoning outcomes, our approach incorporates ground-truth driving actions as hints, guiding the model to produce causal explanations that explicitly link driving decisions to scene context. This structured prompting method significantly reduces nonsensical outputs and minimizes the need for manual correction.
Employing this annotation pipeline, we compile a comprehensive reasoning dataset comprising approximately 45.6k CoT reasoning annotations for the nuPlan dataset and 7.2k annotations for the Waymo E2E dataset. Additionally, we reformat and integrate DriveLM [96], a VQA dataset built on nuScenes and CARLA simulation data, to augment the reasoning data. Additional details and illustrative examples are provided in the Supplementary Material.
@@ 3.4 Reinforcement Fine-tuning
To further improve the performance of AutoVLA and align it with driving requirements and task-specific rewards, we introduce a reinforcement learning-based post-training method. This RFT stage enables the model to perform adaptive reasoning and optimize planning performance. We employ the GRPO algorithm [49], which stabilizes training and improves convergence efficiency. Moreover, the inherent multi-modality of planning, characterized by multiple feasible trajectories in the same scenario, naturally aligns with the group-based optimization framework of GRPO [38].
@@
The final reward function is defined as $r=r_{\text{Driving}}-\lambda_{r}r_{\text{CoT}}$, where the $\lambda_{r}$ denotes the balance weight. The term $r_{\text{Driving}}$ varies across benchmarks. For the nuPlan dataset, we employ the Predictive Driver Model Score (PDMS) [51] as the driving reward, which captures aspects such as safety, comfort, travel efficiency, and other driving quality metrics. For the Waymo E2E dataset, due to the limited availability of Rater Feedback Score (RFS) annotations [52], we use the Average Displacement Error (ADE) as the driving reward. To discourage unnecessarily long reasoning chains, we incorporate a CoT length penalty $r_{\text{CoT}}$ into the reward function. Additional implementation details are provided in the Supplementary Material.
4 Experiments
4.1 Experimental Setup
Datasets. We train the AutoVLA model using a diverse set of real-world and simulation datasets. The nuPlan (Open-Scene) dataset [50, 97] contains 120 hours of large-scale driving data with eight streams of camera data and object annotations. The Waymo end-to-end driving dataset [52] comprises 4,021 20-second driving segments with eight streams of camera views and ego vehicle trajectories, especially focusing on challenging and long-tail scenarios, such as driving through construction areas or risky situations. The nuScenes dataset [53] provides 1,000 urban driving scenes with six camera views. The CARLA-Garage dataset [55] provides over 500,000 frames of camera data from the CARLA simulator. In addition to the collected reasoning data, we utilize the DriveLM dataset [96] for nuScenes and CARLA datasets, by reformatting the VQA pairs to facilitate CoT reasoning.
Benchmarks. We evaluate AutoVLA on both open-loop and closed-loop benchmarks across real-world and simulated environments. Open-loop performance is assessed on two public benchmarks: the NAVSIM benchmark [51] from the nuPlan dataset and the nuScenes benchmark [65]. The NAVSIM benchmark employs PDMS to assess key aspects of driving behavior, such as collision and ego progress. The nuScenes benchmark uses L2 distance and collision rate as evaluation metrics. Additionally, we report our model’s performance on the Waymo end-to-end driving benchmark using the RFS metric, which reflects human-judged planning quality. Closed-loop performance is evaluated on the Bench2Drive benchmark [54] in the CARLA simulator. Bench2Drive contains 44 interactive, closed-loop scenarios under varying locations and weather conditions, using metrics such as success rate, driving score, efficiency, and comfort.
Implementation Details. Each action token corresponds to 0.5 seconds of movement, and the planning horizon is set to 5 seconds. Consequently, the model outputs 10 action tokens, from which a 5-second trajectory can be decoded. For SFT, we use a learning rate of $1\times 10^{-5}$ and the FSDP training strategy. The model is trained for 5 epochs using 8 NVIDIA L40S GPUs. We use a per-GPU batch size of 1 and accumulate gradients over 4 steps, resulting in an effective batch size of 32. The weighting parameters in the SFT loss function are set to $\lambda_{a}=1$ and $\lambda_{cot}=40$. For RFT, we employ the LoRA adapter [98] for parameter-efficient training. The learning rate for RFT is set to $3\times 10^{-5}$, and the KL regularization weight $\beta$ is set to $0.04$. We perform a single policy update at each step, allowing the use of a simplified objective without the need for clipping or tracking the old policy. The model is fine-tuned for $6,000$ steps, and the best-performing checkpoint is selected for evaluation. Additional details are provided in the Supplementary Material.
4.2 Main Results
Data Scaling Results. We train AutoVLA on a mixture of the nuPlan and nuScenes datasets with varying training set sizes (10k, 50k, 100k, and the full 185k samples), with action-only supervision or with additional CoT reasoning supervision. The models are evaluated on the respective standard test sets, and the results are shown in Fig. 4. We observe that increasing the amount of training data consistently improves planning performance on both datasets. In the nuPlan dataset, when using fewer than 50k training samples, CoT reasoning does not outperform action-only in terms of PDMS and Collision Score. This is likely due to the increased difficulty of learning structured reasoning from limited data. However, as the training set size increases, models trained with CoT reasoning surpass those with action-only supervision, highlighting the scalability advantages of reasoning-augmented learning. In the nuScenes dataset, action-only supervision yields better performance in terms of L2 distance and collision rate. This is likely because nuScenes contains mostly simple scenarios that do not require complex reasoning, making CoT training less beneficial in this setting.
@@ 4.2 Main Results
TransFuser [58] 83.88 97.78 92.63 97.97 78.88 92.89 99.98
DRAMA [56] 86.87 98.19 95.18 98.03 81.33 94.17 100.00
Hydra-MDP [57] 91.26 99.07 98.29 95.79 85.20 96.56 100.00
Centaur [59] 92.10 99.23 98.72 96.77 85.96 97.17 99.97
TrajHF [85] 93.95 99.30 97.51 91.72 90.39 98.02 99.81

### v3

@@ 1 Introduction
Autonomous driving systems typically adopt a modular paradigm, decomposing the driving task into different sub-modules, such as perception [li2024bevformer, wang2022detr3d, liang2022bevfusion], prediction [zhou2023qcnext, shi2024mtr++, huang2023gameformer], and planning [huang2024gen, huang2023differentiable, liu2025hybrid]. While this design enables structured development, it may cause error accumulation and a lack of joint optimization across modules, leading to suboptimal performance [shao2023reasonnet, jia2023think]. End-to-end autonomous driving has gained prominence with a unified model architecture that maps raw sensor inputs directly to final driving actions. These models are trained on human driving data, enhancing scalability and human-like behavior. Vision-based approaches have garnered significant interest due to their affordability and ease of deployment [pan2024vlp, Wang_2024_CVPR, hu2022st, fang2023tbp].
@@
However, conventional end-to-end methods [zhou2024vision, cai2024driving, chen2024end, fu2024drive] primarily focus on imitating expert trajectories, lacking essential world knowledge for understanding and reasoning about surrounding environments, particularly in long-tail or challenging scenarios. Recent advances in Vision-Language Models (VLMs) [jaech2024openai, bai2025qwen2, guo2025deepseek] have gained significant interest by introducing models capable of leveraging extensive world knowledge and powerful reasoning. These models have shown strong potential in improving adaptability and scalability across diverse driving scenarios [chen2025automated, li2024driving, xu2024drivegpt4, wang2024omnidrive, ma2024dolphins, winter2025bevdriver, tian2025nuscenes]. Building upon VLMs, Vision-Language-Action (VLA) models extend this capability to action generation, enabling embodied agents, such as robots [kim2024openvla, zhao2025cot, hung2025norasmallopensourcedgeneralist] and autonomous vehicles [arai2025covla, zhou2025opendrivevla], to produce feasible physical actions based on visual observations and language instructions.
@@
Despite recent progress, existing VLA models face two critical limitations in autonomous driving, as illustrated in Fig.˜2. 1) Physically-infeasible or complex structure for action generation. Some models generate textual actions or waypoints directly using VLMs [shao2024lmdrive, renz2025simlingo, hwang2024emma], but these outputs can be physically infeasible and suffer from mode collapse. To address this, recent approaches introduce intermediate meta-actions [jiang2025alphadrive, jiang2024senna, wang2023drivemlm] or latent action tokens [renz2024carllava, fu2025orion, waywe2024lingo], which are then processed by downstream planners or decoders to produce physically feasible trajectories. However, the intermediate representations either break the end-to-end optimization paradigm or increase model complexity and training overhead. 2) Inflexible and inefficient reasoning across diverse scenarios. Most existing models [xing2025openemma, wang2024drivecot] employ a fixed reasoning strategy, lacking the ability to adaptively switch between direct action outputs for straightforward scenarios and chain-of-thought (CoT) reasoning for complex ones. Although DriveVLM [tian2024drivevlm] introduces a dual-process paradigm, it relies on separate modules (i.e., a VLM for slow reasoning and a conventional end-to-end model for fast responses), which results in a complicated architecture, increased training overhead, and limited scalability [ma2025reasoning].
@@
To overcome these limitations, we propose AutoVLA, an end-to-end autonomous driving framework that directly integrates physical action tokens into a pretrained VLM backbone, enabling direct learning of an autoregressive planning policy, as illustrated in Fig.˜1. Our unified architecture seamlessly integrates reasoning and action generation, allowing adaptive switching between direct trajectory generation and CoT reasoning. In supervised fine-tuning (SFT), we leverage both trajectory-only data and CoT reasoning data to equip the model with dual-process capabilities (fast and slow thinking). Furthermore, we propose reinforcement fine-tuning (RFT) [ouyang2022training], utilizing Group Relative Policy Optimization (GRPO) [shao2024deepseekmath] with verifiable planning reward functions. This enables adaptive reasoning that balances planning accuracy and efficiency. The RFT method not only improves planning performance but also runtime efficiency by minimizing unnecessary reasoning.
We extensively evaluate AutoVLA using real-world datasets, including nuPlan [karnchanachari2024towards, dauner2024navsim], Waymo [xu2025wod], nuScenes [caesar2020nuscenes], and simulation datasets such as CARLA [jia2024bench2drive, Jaeger_2023_ICCV]. Experimental results demonstrate that AutoVLA achieves superior performance across various end-to-end autonomous driving benchmarks under both open-loop and closed-loop tests. Empirical results validate that our RFT approach markedly improves planning performance, enables adaptive fast and slow thinking capabilities, and reduces runtime by minimizing redundant reasoning. The main contributions of this paper are summarized as follows:
@@ 2 Related Work
End-to-end Autonomous Driving. End-to-end autonomous driving approaches have made significant advances in recent years [shao2023reasonnet, jia2023think, yuan2024drama, li2024hydra, chitta2022transfuser, sima2025centaur, zhou2024v2xpnp, sun2024sparsedrive, li2024pretrain, zheng2024gaussianad, wang2025generativeaiautonomousdriving]. Methods such as UniAD [hu2023planning] and VAD [jiang2023vad] explicitly integrate multiple driving tasks from perception to planning in a unified Transformer architecture, thereby enhancing planning performance. ParaDrive [weng2024drive] discusses the necessary components within end-to-end driving architectures. Additionally, GenAD [zheng2024genad] and DiffusionDrive [liao2024diffusiondrive] adopt generative models to maintain trajectory continuity and produce multi-modal driving trajectories. However, integrating world knowledge into end-to-end driving systems remains challenging due to bottlenecks in semantic reasoning [zhou2025opendrivevla] and limited adaptability in complex environments [li2024ego].
@@
VLA and VLM for Autonomous Driving. The gap between semantic reasoning and physical actions remains a critical challenge for VLA and VLM in end-to-end autonomous driving. Current research broadly follows three directions. The first directly formulates driving as a language-centric problem, utilizing VLMs for scenario understanding through caption generation [jin2023adapt, park2024vlaad, ding2024hint] or question answering [marcu2024lingoqa, li2024womd]. The second direction leverages VLA or VLM to produce high-level meta-actions or driving decisions [cai2024driving, jiang2025alphadrive, jiang2024senna, wang2023drivemlm], which are used to either supervise [pan2024vlp, xu2024vlm, hegde2025distilling, liu2025vlm] or guide [jiang2024senna, liao2025cot] traditional planners or end-to-end models. Although these approaches facilitate integration, they prevent full end-to-end optimization. Thus, a third direction directly integrates VLMs with action generation into VLA models, enabling the direct prediction of latent action tokens [zhou2025opendrivevla, shao2024lmdrive, renz2025simlingo, waywe2024lingo] or final driving trajectories [hwang2024emma, xing2025openemma, mao2023gpt, zhang2024wisead, tian2024tokenize, qiao2025lightemmalightweightendtoendmultimodal]. However, simple trajectory decoders employed in these methods (e.g., MLP [renz2024carllava, liu2025dsdrivedistillinglargelanguage] or GRU [wang2024drivecot]) may produce impractical trajectories and suffer from modal collapse. To address this issue, ORION [fu2025orion] incorporates generative planners into VLM architectures, enhancing trajectory feasibility but increasing model complexity and computational demands. In our work, we integrate a physical action codebook for vehicle motion into a pretrained VLM to effectively bridge the semantic reasoning and physical action space.
@@
Reinforcement Fine-tuning. RFT [ouyang2022training] has shown considerable promise in enhancing the performance and adaptability of LLMs, as demonstrated in DeepSeek-R1 [guo2025deepseek]. In autonomous driving, Gen-Drive [huang2024gen] and TrajHF [li2025finetuning] employed the RFT to align the trajectory generation model with safety constraints and human driving preferences. RAD [gao2025rad] combined 3D Gaussian splatting to generate scenarios and conduct closed-loop RL training. However, the application of RFT in end-to-end VLM/VLA-based autonomous driving remains nascent. While previous methods, such as AlphaDrive [jiang2025alphadrive], utilize GRPO instead of direct preference optimization (DPO) [rafailov2023direct] to enhance planning performance and ensure training efficiency and stability, they are still limited to simplified settings involving only high-level meta-actions. In this work, we advance this direction by applying RFT to optimize the end-to-end VLA framework in both scene reasoning and low-level planning, and we adopt GRPO to accelerate convergence and enhance training stability.
@@ 3 AutoVLA
The proposed AutoVLA framework consists of two main components, as shown in Fig.˜1. 1) VLM Backbone: It is capable of processing visual and textual input and generating corresponding tokens (reasoning and action), employing a unified autoregressive Transformer decoder. 2) Physical Action Token Generation: We extend the language model decoder to output physical action tokens that directly correspond to vehicle movements. These tokens are designed to comply with physical constraints and can be reliably translated into physically feasible planning trajectories.
Training of AutoVLA is conducted in two stages, as illustrated in Fig.˜3. 1) Supervised Fine-Tuning uses ground-truth trajectory data and distills high-quality reasoning data from a large-scale VLM. 2) Reinforcement Fine-Tuning uses task-specific reward functions to optimize planning performance while improving the running efficiency by minimizing unnecessary reasoning. The details of our model and training process are illustrated below.
@@ 3.1 Framework
Base VLM Model. We adopt Qwen2.5-VL-3B [bai2025qwen2] as the vision-language backbone of AutoVLA. Qwen2.5-VL is a series of powerful multimodal large language models that possess strong visual understanding capabilities, and the open-source nature of the Qwen2.5-VL model facilitates task-specific fine-tuning. The 3B variant offers a good trade-off between efficiency and performance, making it suitable for deployment in onboard devices.
Action Tokenization. To enable trajectory planning within the language model, we discretize continuous vehicle trajectories $\mathbf{P}\in\mathbb{R}^{\tau\times d}$ into a sequence of physical action tokens $\mathbf{a}=[a_{1},\dots,a_{T}]$, where $a_{t}\in\mathcal{A}$, $T$ is the length of the tokenized predicted trajectory and each token is represented by short-term spatial position and heading movement $(\Delta x,\Delta y,\Delta\theta)$. This transforms the planning task into a next-token prediction problem, which can be conducted within the language model. We build our action codebook $\mathcal{A}=\{a_{1},a_{2},\dots,a_{K}\}$ using a K-disk clustering method [zhang2024closed, philion2023trajeglish, wu2024smart], which covers the majority of vehicle movement patterns. Finally, we obtain a vehicle motion codebook that consists of $K=2048$ discrete action tokens. Following [kim2024openvla, pertsch2025fast], these action tokens are incorporated into the VLM as additional tokens (i.e., <action_0>, <action_1>, …). During inference, the model outputs a sequence of these action tokens, which are subsequently decoded into a planning trajectory using the action codebook. More details about action tokenization and trajectory decoding are provided in the Supplementary Material.
Unified Reasoning and Action. AutoVLA unifies reasoning and action generation within a single autoregressive Transformer framework, enabling adaptive switching between fast and slow thinking depending on the driving scenario. In fast thinking mode, AutoVLA directly predicts physical action tokens without generating long CoT reasoning, enabling rapid responses in straightforward scenarios. In contrast, slow thinking mode involves structured CoT reasoning, where the model first analyzes the environment, identifies critical elements, and reasons through potential outcomes before deciding on the final driving action. To enable this dual thinking capability, AutoVLA is trained with a mixture of direct action supervision and reasoning-augmented data. We design system prompts and response formats to support both modes consistently.
@@ 3.2 Reasoning Data
Reasoning data provides high-quality CoT annotations that are essential for training VLMs with reasoning capabilities [fu2025orion]. In driving tasks, reasoning involves understanding complex semantics and interactions in dynamic environments [park2025nuplanqa, wu2025language, qian2024nuscenes, nie2024reason2drive]. Despite its importance, the development of a high-quality, large-scale driving reasoning dataset remains a key challenge due to three major limitations: 1) limited scenario diversity and repetitive examples, 2) inadequate representation of critical perceptual cues, such as traffic signs and vehicle indicator signals, 3) low-quality reasoning process, such as repeatedly stopping at a stop sign without justification.
To address these issues, we propose an automated reasoning annotation pipeline using the advanced Qwen2.5-VL-72B model [bai2025qwen2]. This pipeline enables automatic generation of high-accuracy reasoning annotations and supports knowledge distillation from a large capable model to a more compact target model. The pipeline generates structured reasoning annotations across four key components: detailed scene descriptions, identification of crucial objects, prediction of surrounding agents’ intentions, and determination of appropriate driving actions. To regulate the reasoning outcomes, our approach incorporates ground-truth driving actions as hints, guiding the model to produce causal explanations that explicitly link driving decisions to scene context. This structured prompting method significantly reduces nonsensical outputs and minimizes the need for manual correction.
Employing this annotation pipeline, we compile a comprehensive reasoning dataset comprising approximately 45.6k CoT reasoning annotations for the nuPlan dataset and 7.2k annotations for the Waymo end-to-end driving dataset. Additionally, we reformat and integrate DriveLM [sima2024drivelm], a VQA dataset built on nuScenes and CARLA simulation data, to augment our reasoning data. Additional details and illustrative examples are provided in the Supplementary Material.
@@ 3.4 Reinforcement Fine-tuning
To further improve the performance of AutoVLA and align it with driving requirements and task-specific rewards, we introduce a reinforcement learning-based post-training method. This RFT stage enables the model to perform adaptive reasoning and optimize planning performance. We employ the GRPO algorithm [shao2024deepseekmath], which stabilizes training and improves convergence efficiency. Moreover, the inherent multi-modality of planning, characterized by multiple feasible trajectories in the same scenario, naturally aligns with the group-based optimization framework of GRPO [jiang2025alphadrive].
@@
The final reward function is defined as $r=r_{\text{Driving}}-\lambda_{r}r_{\text{CoT}}$, where the $\lambda_{r}$ denotes the balance weight. The term $r_{\text{Driving}}$ varies across benchmarks. For the nuPlan dataset, we employ the Predictive Driver Model Score (PDMS) [dauner2024navsim] as the driving reward, which captures aspects such as safety, comfort, travel efficiency, and other driving quality metrics. For the Waymo end-to-end driving dataset, due to the limited availability of Rater Feedback Score (RFS) annotations [xu2025wod], we use the Average Displacement Error (ADE) as the driving reward. To discourage unnecessary long reasoning chains, we incorporate a CoT length penalty $r_{\text{CoT}}$ into the reward function. Additional implementation details are provided in the Supplementary Material.
@@
TransFuser [chitta2022transfuser] 83.88 97.78 92.63 97.97 78.88 92.89 99.98
DRAMA [yuan2024drama] 86.87 98.19 95.18 98.03 81.33 94.17 100.00
Hydra-MDP [li2024hydra] 91.26 99.07 98.29 95.79 85.20 96.56 100.00
Centaur [sima2025centaur] 92.10 99.23 98.72 96.77 85.96 97.17 99.97
TrajHF [li2025finetuning] 93.95 99.30 97.51 91.72 90.39 98.02 99.81
@@
4 Experiments
4.1 Experimental Setup
Datasets. We train the AutoVLA model using a diverse set of real-world and simulation datasets. The nuPlan (Open-Scene) dataset [karnchanachari2024towards, openscene2023] contains 120 hours of large-scale driving data with eight streams of camera data and object annotations. The Waymo end-to-end driving dataset (Waymo E2E) [xu2025wod] comprises 4,021 20-second driving segments with eight streams of camera views and ego vehicle trajectories, especially focusing on challenging and long-tail scenarios, such as driving through construction areas or risky situations. The nuScenes dataset [caesar2020nuscenes] provides 1,000 urban driving scenes with six camera views. The CARLA-Garage dataset [Jaeger_2023_ICCV] provides over 500,000 frames of simulation camera data. In addition to the collected reasoning data, we utilize the DriveLM dataset [sima2024drivelm] for nuScenes and CARLA datasets, by reformatting the VQA pairs to facilitate CoT reasoning.
Benchmarks. We evaluate AutoVLA on both open-loop and closed-loop benchmarks across real-world and simulated environments. Open-loop performance is assessed on two public benchmarks: the NAVSIM benchmark [dauner2024navsim] from the nuPlan dataset and the nuScenes benchmark [hu2023planning]. The NAVSIM benchmark employs PDMS to assess key aspects of driving behavior, such as collision and ego progress. The nuScenes benchmark uses L2 distance and collision rate as evaluation metrics. Additionally, we report our model’s performance on the Waymo end-to-end driving benchmark using the RFS metric, which reflects human-judged planning quality. Closed-loop performance is evaluated on the Bench2Drive benchmark [jia2024bench2drive] in the CARLA simulator. Bench2Drive contains 44 interactive, closed-loop scenarios under varying locations and weather conditions, using metrics such as success rate, driving score, efficiency, and comfort.
Implementation Details. Each action token corresponds to 0.5 seconds of movement, and the planning horizon is set to 5 seconds. Consequently, the model outputs 10 action tokens, from which a 5-second trajectory can be decoded. For SFT, we use a learning rate of $1\times 10^{-5}$ and the FSDP training strategy. The model is trained for 5 epochs using 8 NVIDIA L40S GPUs. We use a per-GPU batch size of 1 and accumulate gradients over 4 steps, resulting in an effective batch size of 32. The weighting parameters in the SFT loss function are set to $\lambda_{a}=1$ and $\lambda_{cot}=40$. For RFT, we employ the LoRA adapter [hu2022lora] for parameter-efficient training. The learning rate for RFT is set to $3\times 10^{-5}$, and the KL regularization weight $\beta$ is set to $0.04$. We perform a single policy update at each step, allowing the use of a simplified objective without the need for clipping or tracking the old policy. The model is fine-tuned for $6,000$ steps, and the best-performing checkpoint is selected for evaluation. Additional details are provided in the Supplementary Material.
4.2 Main Results
Data Scaling Results. We train AutoVLA on a mixture of the nuPlan and nuScenes datasets with varying training set sizes (10k, 50k, 100k, and the full 185k samples), with action-only supervision or with additional CoT reasoning supervision. The models are evaluated on the respective standard test sets, and the results are shown in Fig.˜4. We observe that increasing the amount of training data consistently improves planning performance on both datasets. In the nuPlan dataset, when using fewer than 50k training samples, CoT reasoning does not outperform action-only in terms of PDMS and Collision Score. This is likely due to the increased difficulty of learning structured reasoning from limited data. However, as the training set size increases, models trained with CoT reasoning surpass those with action-only supervision, highlighting the scalability advantages of reasoning-augmented learning. A similar trend is observed on the nuScenes dataset: as the training set size increases, models trained with CoT reasoning consistently outperform those trained with action-only data in terms of L2 distance and collision rate.
Table 2: Runtime Analysis of Fast & Slow Thinking Modes in AutoVLA
Thinking Mode Min. (s) Max. (s) Avg. (s)
Fast Thinking 0.997 1.116 1.072
Slow Thinking 7.607 13.706 10.518